OpenStudy (anonymous):

Paul is a cab driver and he charges a base fee of $8.80 for every fare, plus $1.20 for every mile traveled. If n represents the number of miles traveled, which of the following sentences can be used to find C, the amount Paul charges for each fare? C = 8.8 - 1.2n C = 8.8n + 1.2 C = 8.8 + 1.2n C = 8.8 × 1.2n

OpenStudy (anonymous):

C = 8.8 + 1.2n

OpenStudy (anonymous):

The height of a ball that was batted into the air at 160 feet per second is a function of t, the time in seconds after the ball was hit. The height is determined by subtracting 16 times the square of t from 160 times t. Which equation can be used to find t when the ball is 300 feet high? 160(t2 - t) = 300 160t2 - 16t = 300 (160 - 16)t2 = 300 160t - 16t2 = 300

OpenStudy (anonymous):

160t - 16t2 = 300 should do it.

OpenStudy (anonymous):

To avoid a large, shallow reef, a ship set a course from point A and traveled 20 miles east to point B. The ship then turned and traveled 35 miles south to point C. If the ship could have traveled in a straight line from point A to point C, about how many miles could it have saved? 19 miles 15 miles 40 miles 5 miles

OpenStudy (anonymous):

15 Miles could've been saved.